# AL-Bench: A Benchmark for Automatic Logging

### v2

@@ 3. AL-Bench
In this section, we introduce our benchmark, AL-Bench, which builds on insights from previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023). AL-Bench consists of two key parts: static evaluation and dynamic evaluation, both of which we will introduce in detail. The static evaluation component includes a high-quality dataset with 42,224 instances and five metrics, based on the static evaluation method outlined in Section 2.2.2. The dynamic evaluation component introduces a separate dataset with 2,238 instances and four metrics, following the dynamic evaluation approach, which will be described later in this section.
@@ 3.1. Static Evaluation
3.1.1. Static Evaluation Method.
Static evaluation focuses on log statement components—position, verbosity, and message—and has been the primary method in prior studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), with details provided in Section 2.2.2.
3.1.2. Dataset Construction
As discussed in Section 2.2.1, the quality of the evaluation dataset is crucial for assessing the performance of tools. The strictest existing rules for dataset collection—500 commits, 10 contributors, and 10 stars—are inadequate to ensure the quality of the evaluation dataset. Drawing inspiration from the use of GitHub repository stars as an effective metric for identifying high-quality code datasets  (Jiang et al., 2024), and considering the unique characteristics of log statements, we propose three critical criteria for dataset selection: repositories must have at least 10,000 stars, 1,000 log statements, and 500 log-related issues to qualify for inclusion in our dataset. These standards guarantee well-maintained, widely used projects where developers prioritize high-quality log statements. Given the diverse requirements for log statements in different scenarios, we ultimately selected 10 projects based on domain diversity, industrial applications, and their relevance to prior logging studies (Chen and Jiang, 2017, 2019; Li et al., 2018, 2019; Ding et al., 2022, 2023; Li et al., 2020). As shown in Table 1, our final dataset includes projects with a total of 22,787 code snippets and 42,224 log statements, covering a wide range of logging needs and practices. The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms, ensuring diversity in logging scenarios. Each domain presents unique requirements for log statements.
For example, database management systems such as DBeaver and Doris prioritize minimizing the impact of logging on high performance. Task scheduling systems, including DolphinScheduler, rely on logging to trace task dependencies and monitor runtime statistics. Similarly, distributed systems like Hadoop and Zookeeper require robust logging practices to address challenges in distributed coordination, fault tolerance, and scalability, which differ from the requirements of other command logs. Messaging systems such as Kafka and Pulsar have adapted logging practices to trace message flows, ensure reliable message delivery, and debug asynchronous communication. Meanwhile, IoT platforms like ThingsBoard utilize logging to manage device connectivity, monitor data streams, and enable real-time system oversight. Identity and access management systems such as Keycloak prioritize protecting sensitive information in logs to prevent privacy breaches. By including diverse projects across these domains, the dataset ensures comprehensive coverage of different logging practices, making it suitable for benchmarking and analyzing log-related tasks under various operational contexts.
In addition to emphasizing data quality, we also addressed the potential risk of data contamination. Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk. Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data. Furthermore, we wrapped the code snippets in a common class and standardized the formatting using Google-Java-Format (Google, [n. d.]). This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure. These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024). Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024). However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation. In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current. After completing the necessary appeal actions, we finalized our static evaluation dataset.
Table 1. Details of AL-Bench dataset.
Dataset Domain Code snippets with log statements Total log statements
Dbeaver Database Management 1,182 1,725
Dolphinscheduler Task Scheduling 898 1,918
Doris High Performance Database 1,686 2,965
Flink Data Processing 1,956 3,268
Hadoop Distributed Storage 8,751 16,002
Kafka Messaging Systems 1,706 3,421
Keycloak Identity and Access Management 533 1,004
Pulsar Messaging Systems 3,590 7,254
Thingsboard IoT Platform 1,601 2,731
Zookeeper Distributed Coordination 884 1,936
Total - 22,787 42,224
@@ 3.1.3. Metrics
Building on previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), We adopted five metrics for static evaluation. In addition to the previously established metrics, we propose two new ones: Dynamic Variable Accuracy and Static Text BLEU, designed to better reflect the real-world quality requirements of log statements.
Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements. Some common levels, like ”info,” refer to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. Each level refers to different meanings so for Level Accuracy we strictly compared the level of predictions and source code and used the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
@@
Metric 2: Position Accuracy (PA): Position Accuracy focuses on the precise location of log statements within the source code, which is crucial for tracking and debugging software behavior. Correct placement of log statements helps in accurately tracing the execution flow and diagnosing issues. For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code. This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
@@
Metric 3: Message Accuracy (MA): Message Accuracy evaluates how accurately the predicted log messages match the ground truth log messages, which is essential for providing meaningful and relevant information during runtime. The content of log messages helps developers understand the system’s behavior, and inaccuracies in message generation can lead to confusion or missed insights during debugging. For Message Accuracy, we compare the predicted log messages to the actual messages in the source code. This metric is calculated by determining the number of log messages that are fully identical to the ground truth $M_{c}$ and dividing it by the total number of log messages $N_{a}$, yielding the accuracy value: $MA=\frac{M_{c}}{N_{a}}$.
Metric 4: Dynamic Variable Accuracy (DVA): Dynamic Variable Accuracy focuses on the dynamic variable in log message, for example, (”The server is running, {}”, status), in this message, status is regarded as the dynamic variable in log message which might vary according to the runtime behaviors of software. And in the log message (”The server run on the ports, {}”, args.status ? localPort : remotePort), the whole expression (args.status ? localPort : remotePort) is reckoned as the dynamic variable which could decide the port information recorded in log files. We aim to use this metric to ensure that the dynamic information recorded in logs remains consistent. We extract the dynamic variables from both the source code and predictions and then compare them to record instances where they match exactly. This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DVA=\frac{DP_{C}}{N_{a}}$.
@@
Metric 5: Static Text BLEU (STB): Static Text BLEU focuses on the static part of the log message. Unlike the dynamic variable, the static part always records the same information in log files, which will not vary due to the runtime behavior of software. For example, in the log message (”The server is running, {}”, status), ”The server is running, {}” is regarded as the static part. Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) metric to evaluate the quality of the static text. In our implementation, we use the BLUE-DM variant (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method.
@@ 3.2. Dynamic Evaluation
Figure 6. The general workflow of dynamic evaluation. First, compile the project and run the unit test to obtain ground truth logs. Then, replace log statements with predictions, rerun the test to generate new logs, and finally analyze the results.
3.2.1. Dynamic Evaluation Method
Different from static evaluation, dynamic evaluation focuses on compiling the code and runtime-generated logs, addressing static evaluation’s inability to verify code compilability and runtime logs. To directly assess runtime logs, we generate them using unit tests, which are widely used in software development to verify code functionality in isolated scenarios. Unit tests are readily available in most projects and they are designed to test the functionality and behaviors of code when facing different situations. They offer a natural method for simulating realistic situations, allowing generating logs without the need for complex runtime environments. Figure 6 demonstrates the general workflow of dynamic evaluation. The process begins with compiling the source code and executing the unit tests to obtain logs from the original log statements. Using tools like Jacoco (Jacoco, [n. d.]) and SureFire (surefire, [n. d.]), we collect the logs and remove the log statements covered by unit tests in the source code. Next, we input the modified source code into an automatic logging tool to generate predicted log statements. Then we inserted the predicted log statements back into the source code, replacing the original log statements. After recompiling the modified code, we rerun the unit tests to capture the logs produced by the inserted predicted log statements. The whole process provides us with two sets of logs—those generated by the original log statements and those generated by the predicted log statements. Finally, we evaluate the effectiveness of the predicted log statements using two key metrics: Compilation Success Rate and Log File Similarity. Compilation Success Rate ensures that the predicted log statements do not introduce compilation errors, while Log File Similarity measures the similarity between the logs generated by the predicted log statements and those generated by the original log statements. In the following sections, we will detail how we built the dataset for dynamic evaluation and introduce the specific metrics used to measure the performance of the automatic logging tools.
3.2.2. Dataset Construction
@@
To build the dynamic evaluation dataset, we begin by compiling the entire project to ensure all dependencies are resolved and the project is ready for execution. Next, we systematically identify all available unit tests within the project. For each unit test, we execute it individually while employing the Jacoco Plugin (Jacoco, [n. d.]) to trace code coverage, specifically identifying whether the unit test interacts with or covers any log statements in the codebase. Simultaneously, we use the SureFire Plugin (surefire, [n. d.]) to capture the logs generated during the execution of the unit tests.
By correlating Jacoco coverage data with SureFire logs, we can match specific code snippets containing log statements to the corresponding unit tests that cover them, along with the runtime logs they generate. This process enables us to construct a comprehensive dataset consisting of triples: the code snippet, the unit test that triggers it, and the recorded logs. These triples are critical for dynamic evaluation, as they provide the ground truth for assessing the quality and effectiveness of predicted log statements in actual runtime scenarios.
To construct a dynamic evaluation dataset, it is essential to select high-quality projects with comprehensive unit tests, as these tests provide a realistic simulation of diverse production environments. Each instance of dynamic evaluation requires recompiling the project and executing the test to collect logs, making the process highly time-consuming. To balance this intensive time requirement with the need for sufficient dataset diversity and quantity, we employed Hadoop as our dynamic evaluation platform. This approach allowed us to build a dataset of 2,238 instances that balances diversity with sufficient size. Additionally, we open-source the entire suite of tools used in this evaluation process, enabling researchers and organizations to easily deploy and customize their own dynamic evaluation datasets.
@@ 3.2.3. Metrics
In dynamic evaluation, we proposed four metrics to assess the performance of logging tools: Compilation Success Rate, Log Similarity, False Positive Log Generation Rate, and False Negative Log Generation Rate, which we will introduce below.
Metric 6: Compilation Success Rate (CSR): Compilation Success Rate measures the syntactic correctness of the predicted log statements. As discussed earlier, after replacing the original log statements with the predictions, we recompile the project. However, due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all predictions can be successfully compiled. We recorded the successfully compiled code snippet number as $C_{s}$ and all code snippets as $C_{a}$. The metric is then calculated as: $CSR=\frac{C_{s}}{C_{a}}$.
Metric 7: Log Similarity Metrics Group (Cosine Similarity, BLEU, ROUGE): This metric evaluates how closely logs generated by predicted statements match those produced by ground truth statements. To eliminate unnecessary differences, we remove log headers (e.g., timestamps), retaining only log content. For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004). Cosine Similarity, commonly used in text analysis, calculates the cosine of the angle between two TF-IDF (Jones, 2021) vectors, yielding 1 for identical vectors and 0 for orthogonal ones. Using TF-IDF, we down-weight frequent terms, emphasizing distinctive content in logs. This method effectively captures the similarity between meaningful log content, filtering out redundant information for a more accurate relevance measure. ROUGE, on the other hand, focuses on recall by comparing n-grams between the predicted and reference logs. It evaluates how much of the reference content is preserved in the prediction. The most commonly used variant is ROUGE-N, which calculates the overlap of n-grams between two texts.
Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. the metric is calculated as: $FRLR=\frac{FP}{P}$.
Metric 9: False Negative Log Generation Rate (FNLR): This metric evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs. It highlights instances where the predicted logs miss important events or information. The number of false negative instances is recorded as $FN$, and the total number of predictions is $P$. The metric is calculated as: $FNLR=\frac{FN}{P}$.

### v3

@@ 3. AL-Bench
AL-Bench is a novel benchmark designed to evaluate automatic logging tools in codebases systematically. The benchmark comprises 21,804 code snippets and 39,600 log statements extracted from 10 most popular open-source repositories, selected for their active maintenance, diverse application domains, and representation of modern software engineering practices. AL-Bench employs a dual evaluation framework: Static Evaluation evaluates the similarity of log statements (e.g., position, level, variable, message) between predicted and reference log statements, while Dynamic Evaluation assesses code compileability and the similarity of logs printed by those log statements during runtime execution. Both evaluation methods will be described later in this section.
3.1. Dataset Construction
As discussed in Sec. 1, the quality of the evaluation dataset is crucial for assessing the performance of tools. Building on empirical evidence that GitHub repository stars correlate with code quality (Jiang et al., 2024), we determine dataset inclusion criteria for log statement analysis by requiring repositories to have $\geq$ 10,000 stars and $\geq$ 500 log-related issues to ensure community validation and logging relevance. Candidate repositories are then ranked by total log statement count to prioritize those with extensive logging practices. From the top of this ranked list, we manually curate a final dataset of 10 repositories to ensure that the set covers different priorities in the logging effectiveness-efficiency trade-off, including performance-critical systems (requiring minimal logs) and trace-intensive applications (requiring detailed logs). After that, we ensure that all of these repositories are actively maintained, with frequent commits (¿50/month), a high issue resolution rate (¿70%), and recent updates (within the past 6 months). This dual-phase approach balances the quantitative scale with domain representativeness.
As shown in Table 1, our final dataset includes projects with a total of 21,804 code snippets and 39,600 log statements, covering a wide range of logging needs and practices. The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms. These repositories have distinct logging objectives (targeted downstream tasks) based on the specific needs of their respective scenarios and exhibit unique effectiveness-efficiency trade-offs (more logging vs less logging), as shown in Sec. 2.2. For example, database systems like DBeaver minimize logging overhead to sustain performance, while task schedulers such as DolphinScheduler emphasize dependency-tracing logs for runtime monitoring. Distributed systems (e.g., Hadoop) enforce rigorous logging for fault tolerance and distributed coordination, contrasting with messaging platforms like Kafka, which track message flows to ensure delivery reliability. IoT solutions (e.g., ThingsBoard) leverage logs for real-time device connectivity oversight, whereas security-focused systems like Keycloak prioritize log anonymization to safeguard sensitive data.
These diverse logging requirements not only demonstrate the heterogeneity and inherent contradictions in logging practices (e.g., performance-sensitive scenarios demand minimal logging overhead, while debugging scenarios necessitate comprehensive event records) but also substantiate the impossibility of establishing universal standards for assessing log code quality. By documenting real-world logging practices from community-vetted, widely-used projects across domains, we establish a foundation for evaluating logging strategies in their native environments. This domain-rooted dataset supports comprehensive analysis of automated logging techniques across diverse operational scenarios.
Table 1. Details of AL-Bench dataset.
Dataset Domain # Code w/ LogStmt # LogStmt
Dbeaver Database Management 1,178 1,707
Dolphinscheduler Task Scheduling 883 1,855
Doris High Performance Database 1,635 2,860
Flink Data Processing 1,916 3,150
Hadoop Distributed Storage 8,351 14,987
Kafka Messaging Systems 1,667 3,308
Keycloak Identity and Access Management 491 926
Pulsar Messaging Systems 3,279 6,347
Thingsboard IoT Platform 1,544 2,588
Zookeeper Distributed Coordination 860 1,872
Total - 21,804 39,600
@@ 3.2. Static Evaluation
3.2.1. Task Formulation
We structure each entry in our dataset as a tuple $<Code_{w/o~{}LogStmt},LogPos,LogStmt>$. As illustrated in Figure 3, $Code_{w/o~{}LogStmt}$ represents the input code context with one log statement deliberately removed, $LogPos$ indicates the ground truth position for the missing log statement, and $LogStmt$ represents the exact log statement to be predicted. This process yielded 39,600 high-quality instances, providing a robust foundation for evaluating log generation models.
@@ 3.2.2. Metrics
We introduce six static evaluation metrics to assess the similarity between the generated and reference log statements.
@@
Metric 1: Position Accuracy (PA): PA focuses on the precise line of log statements within the source code. The correct placement of log statements helps accurately trace the execution flow and diagnose issues. For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code. This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
Metric 2: Level Accuracy (LA): LA evaluates the exact match between predicted and reference log levels, which are essential for prioritizing operational events in DevOps pipelines. Common log levels carry distinct semantic implications, like ”info,” referring to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. LA is calculated as the ratio of correctly predicted levels $L_{c}$ to the total log statements $LA=\frac{L_{c}}{N_{a}}$.
Metric 3: Average Level Distance (ALD): ALD further quantifies the severity deviation of mispredicted levels from the reference level. We assign ordinal values to log levels: trace: 0, debug: 1, info: 2, warn: 3, error:4, fatal: 5, and compute the absolute difference between predicted level $L_{p}^{(i)}$ and oracle level $L_{a}^{(i)}$ for each log. ALD is the mean deviation across all instances: $ALD=\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}\left|L_{p}^{(i)}-L_{a}^{(i)}\right|$.
@@
Metric 4: Message Accuracy (MA): MA evaluates how accurately the predicted log messages match the oracle, which is essential for providing meaningful and relevant information during runtime. The content of log messages helps developers understand the system’s behavior, and inaccuracies in message generation can lead to confusion or missed insights during debugging. For Message Accuracy, we compare the predicted log messages to the actual messages in the source code. This metric is calculated by determining the number of log messages that are fully identical to the ground truth $M_{c}$ and dividing it by the total number of log messages $N_{a}$, yielding the accuracy value: $MA=\frac{M_{c}}{N_{a}}$.
Metric 5: Dynamic Expression Accuracy (DEA): DEA evaluates whether generated logs preserve the structural integrity of runtime expressions, including both individual variables and composite logic (e.g., ternary operators, arithmetic). For example, in the log template:(”The server run on the ports, {}”, args.status ? localPort : remotePort), the conditional expression (args.status ? localPort : remotePort) is treated as a single semantic unit. We aim to use this metric to ensure that the dynamic information recorded in logs remains consistent. This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DEA=\frac{DP_{C}}{N_{a}}$.
@@
Metric 6: Static Text Similarity (STS) (with BLEU or ROUGE): STS focuses on the static part of the log message. Unlike the dynamic variable, the static part always records the same information in log files, which will not vary due to the runtime behavior of software. For example, in the log message (”The server is running, {}”, status), ”The server is running, {}” is regarded as the static part. Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) metrics to evaluate the quality of the static text. In our implementation, we use the DM variant of BLEU (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method — coupled with ROUGE-L to holistically assess both lexical precision and long-sequence coherence. Specifically, ROUGE-L focuses on the longest common subsequence that effectively captures key operational patterns (e.g., error codes or API call chains) in multi-line logs, while BLEU’s n-gram overlap measurement complements it by evaluating template fidelity at the token level.
Figure 3. $Code_{w/o\ \ LogStmt}$ indicates the code without one log statement, the $LogPos$ means the position of this log statement, $LogStmt$ is the log statement itself. Those three structure the evaluation tuple.
@@ 3.3. Dynamic Evaluation
Figure 4. The general workflow of dynamic evaluation. First, compile the project and run the unit test to obtain ground truth logs. Then, replace log statements with predictions, re-run the test to generate new logs, and finally analyze the results.
3.3.1. Task Formulation
In dynamic evaluation, we provide the runtime perspective of evaluating the automatic logging tools. Two key evaluation metrics are employed: (1) compilability—ensuring the code with predicted log statements compiles without errors, and (2) log file similarity—measuring the alignment between log files generated by the predicted log statements and the oracle through textual similarity analysis. These criteria jointly validate both the functional correctness of log integration and the relevance of logged content.
As illustrated in Figure 4, the core innovation of our methodology lies in utilizing unit tests to emulate runtime log generation under controlled conditions. By executing these tests, we synthetically replicate software runtime states with minimal computational overhead, thereby enabling systematic collection of runtime logs that mirror real-world execution patterns for downstream analysis. We define task formulation as a tuple $<Code_{w/o\ \ LogStmt},Code_{w/\ \ LogStmt},Logs>$, where $Code_{w/o\ \ LogStmt}$ is the code snippets without log statements covered by unit test, $Code_{w/\ \ LogStmt}$ is the code snippets with predicted log statement, and $Logs$ is the logs produced by those log statements in $Code_{w/\ \ LogStmt}$ via unit tests. If the $Code_{w/\ \ LogStmt}$ is not compiliable, $Logs$ is null.
@@ 3.3.1. Task Formulation
To build the dynamic evaluation pair based on our dataset, we begin by compiling the project to ensure all dependencies are resolved and the project is ready for execution. Next, we systematically identify all available unit tests within the project. For each unit test, we execute it individually while employing the Jacoco Plugin (Jacoco, [n. d.]) to trace code coverage, specifically identifying whether the unit test covers any log statements in the codebase. Simultaneously, we use the SureFire Plugin (surefire, [n. d.]) to capture the logs generated during the execution of the unit tests. By correlating Jacoco coverage data with SureFire logs, we can match specific code snippets containing log statements to the corresponding unit tests that cover them, along with the runtime logs they generate. We finally built 2,238 instances for dynamic evaluation.
@@ 3.3.2. Metrics
We propose four dynamic evaluation metrics to assess the similarity of logs printed by the generated and reference log statements in runtime execution.
Metric 7: Compilation Success Rate (CSR): CSR measures the compilability of the predicted log statements. Due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all code snippets with predicted log statements can successfully compile. We recorded the successfully compiled code snippet number as $C_{s}$ and the evaluation pair number as $C_{a}$. The metric is then calculated as: $CSR=\frac{C_{s}}{C_{a}}$.
Metric 8: Log File Similarity (LFS) (with COS, BLEU, ROUGE): LFS evaluates how closely logs generated by predicted statements match those produced by ground truth statements. We remove log headers (e.g., timestamps) and only assess the main message body of each log entry to eliminate unnecessary differences. For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (COS) (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004). Cosine Similarity, commonly used in text analysis, calculates the cosine of the angle between two TF-IDF (Jones, 2021) vectors, yielding 1 for identical vectors and 0 for orthogonal ones. Using TF-IDF, we down-weight frequent terms, emphasizing distinctive content in logs. This method effectively captures the similarity between meaningful log content, filtering out redundant information for a more accurate relevance measure. ROUGE, on the other hand, focuses on recall by comparing n-grams between the predicted and reference logs. It evaluates how much of the reference content is preserved in the prediction. The most commonly used variant is ROUGE-N, which calculates the overlap of n-grams between two texts.
Metric 9: False Positive Log Generation Rate (FPLR): FPLR measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. the metric is calculated as: $FRLR=\frac{FP}{P}$.
Metric 10: False Negative Log Generation Rate (FNLR): FNLR evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs. It highlights instances where the predicted logs miss important events or information. The number of false negative instances is recorded as $FN$, and the total number of predictions is $P$. The metric is calculated as: $FNLR=\frac{FN}{P}$.